OpenStudy (anonymous):

solve x^2-9x+20=0 solve x^2-9x+20=0 @Mathematics

myininaya (myininaya):

20=-5(-4) -9=-5-4

OpenStudy (mimi_x3):

(x-5)(x-4) = 0 x = 5 , 4

myininaya (myininaya):

(x-5)(x-4)=0